OpenStudy (anonymous):

How Can I Solve This ?? f(x) = 4x - 3 / 7 f(0) Pls.Help me :DDD

OpenStudy (anonymous):

f(0) means substitute 0 where you have x in the equation f(x) So: f(0)=4*(0)-3/7

OpenStudy (anonymous):

f(0)= 4(0)-(3/7) f(0)=-3/7

OpenStudy (anonymous):

is it okay to have a negative numerator ??

OpenStudy (anonymous):

Sure, why not.

OpenStudy (anonymous):

okay thanks :)))